# Beyond the Limits: A Survey of Techniques to Extend the Context Length in Large Language Models

### v2

@@ 6 Hardware-aware Transformers
IO-awareness and Resource Management
@@
This IO-aware approach is not exclusive to attention-based transformers; similar techniques have been applied to expedite SSMs. SSMs, emerging as alternatives to transformers due to linear scalability and convolution implementation feasibility, present challenges in convolution-dominated computation time during training. FlashConv Fu et al. (2022) addresses this by leveraging the Cooley-Tukey decomposition of the FFT into a series of diagonal matrix multiplication, to take advantage of fast tensor cores. To accommodate longer sequences, FlashConv utilizes the recurrent properties of SSMs, allowing convolution to occur in different portions sequentially. Mamba further enhances SSMs through innovative techniques such as kernel fusion, parallel scan, and recomputation, leveraging modern accelerators like GPUs for efficient memory hierarchy utilization Gu and Dao (2023).
Multi-device Distributed Attention
Both FlashAttention and BPT leverage distinct streaming multiprocessors (SMs) in GPUs for parallel processing of different blocks. However, the limited size of SRAM imposes constraints on sequence length. Encouragingly, this concept can be expanded to accommodate very long sequences by distributing attention computation across multiple GPUs, as proposed by Ring Attention Liu et al. (2023a). This innovative approach enables block-wise self-attention computation, seamlessly overlapping communication of key-value blocks among devices with the computation of each block on devices. As a result, it facilitates the processing of sequences several times longer than BPT, showcasing scalability across the device count. Another example of distributing attention computation across multiple devices is demonstrated by LongNet Ding et al. (2023). LongNet possesses the capability to compute multiple attentions, each with a distinct dilated sparsity pattern. These computations operate independently and can be distributed over multiple devices, with each device corresponding to a single dilated pattern. This collective approach facilitates the processing of longer sequences.
Memory Management
@@
Another notable memory management technique is LLM-in-Flash  Alizadeh et al. (2023), which leverages the larger size of flash memory compared to Dynamic Random Access Memory (DRAM). This method runs an LLM during inference efficiently by storing model parameters in flash memory and bringing them to DRAM on demand. To balance the lower bandwidth of Flash memory, the paper introduces an inference cost model considering flash memory characteristics. The technique incorporates sparsity awareness in feedforward layers and context-adaptive loading of the model. Although not specifically used to increase sequence length, this method has the potential to load a model up to twice the size of the available DRAM. This capacity could be harnessed to handle longer sequences while ensuring practical data transfer between DRAM and Flash.
@@ 7 Conclusion and Future Directions
In this survey, a systematic review of different approaches for efficiently extending the sequence length in LLMs is provided. We start with the motivation of the work and the necessity of handling long sequences by LLMs. We then discuss the methods that encompass architectural changes, such as positional encoding modification, and attention mechanism modification,designed to improve long sequences handling without significantly increasing the computational cost. We further delve into the methods that can be applied to different phases, such as training, fine-tuning and inference, which efficiently improve the LLM’s capability of processing long sequences. These techniques not only address the immediate challenges posed by sequence length limitations but also pave the way for more complex and contextually aware LLMs.

### v3

@@ 6 Hardware-aware Transformers
IO-awareness and Resource Management.
@@
This IO-aware approach is not exclusive to attention-based transformers; similar techniques have been applied to expedite SSMs. SSMs, emerging as alternatives to transformers due to linear scalability and convolution implementation feasibility, present challenges in convolution-dominated computation time during training. FlashConv Fu et al. (2023) addresses this by leveraging the Cooley-Tukey decomposition of the FFT into a series of diagonal matrix multiplication, to take advantage of fast tensor cores. To accommodate longer sequences, FlashConv utilizes the recurrent properties of SSMs, allowing convolution to occur in different portions sequentially. Mamba Gu and Dao (2023) further enhances SSMs through innovative techniques such as kernel fusion, parallel scan, and recomputation, leveraging modern accelerators like GPUs for efficient memory hierarchy utilization.
Multi-device Distributed Attention.
Both FlashAttention and BPT leverage distinct streaming multiprocessors in GPUs for parallel processing of different blocks. However, the limited size of SRAM imposes constraints on sequence length. Encouragingly, this concept can be expanded to accommodate very long sequences by distributing attention computation across multiple GPUs, as proposed by Ring Attention Liu et al. (2023a). This innovative approach enables block-wise self-attention computation, seamlessly overlapping communication of key-value blocks among devices with the computation of each block on devices. As a result, it facilitates the processing of sequences several times longer than BPT, showcasing scalability across the device count. Another example of distributing attention computation across multiple devices is demonstrated by LongNet Ding et al. (2023). LongNet possesses the capability to compute multiple attentions, each with a distinct dilated sparsity pattern. These computations operate independently and can be distributed over multiple devices, with each device corresponding to a single dilated pattern. This collective approach facilitates the processing of longer sequences.
Memory Management.
@@
Another notable memory management technique is LLM-in-Flash  Alizadeh et al. (2023), which leverages the larger size of flash memory compared to Dynamic Random Access Memory (DRAM). This method runs an LLM during inference efficiently by storing model parameters in flash memory and bringing them to DRAM on demand. To balance the lower bandwidth of Flash memory, the paper introduces an inference cost model considering flash memory characteristics. The technique incorporates sparsity awareness in feedforward layers and context-adaptive loading of the model. Although not specifically used to increase sequence length, this method has the potential to load a model up to twice the size of the available DRAM. This capacity could be harnessed to handle longer sequences while ensuring practical data transfer between DRAM and flash memory.
@@ 7 Conclusion and Future Directions
In this survey, a systematic review of different approaches for efficiently extending the sequence length in LLMs is provided. We start with the motivation of the work and the necessity of handling long sequences by LLMs. We then discuss the methods that encompass architectural changes, such as positional encoding modification, and attention mechanism modification,designed to improve long sequences handling without significantly increasing the computational cost. We further explore the methods that can be applied to different phases, such as training, fine-tuning and inference, which efficiently improve the LLM’s capability of processing long sequences. These techniques not only address the immediate challenges posed by sequence length limitations but also pave the way for more complex and contextually aware LLMs.